# BAD-NEUS: Rapidly converging trajectory stratification

### v1

@@ V Two-dimensional test system
$V(u,v)=\frac{1}{20}\sum_{i=1}^{4}C_{i}\exp[a_{i}(u-v_{i})^{2}\\ +b_{i}(u-u_{i})(v-v_{i})+c_{i}(v-v_{i})^{2}].$ (34)
For all results shown, we use $C_{i}=\{-200,-100,-170,15\}$, $a_{i}=\{-1,-1,-6.5,0.7\}$, $b_{i}=\{0,0,11,0.6\}$, $c_{i}=\{-10,-10,-6.5,0.7\}$, $u_{i}=\{1,-0.27,-0.5,-1\}$, $v_{i}=\{0,0.5,1.5,1\}$. The potential with these parameter choices is shown in Figure 1. The presence of multiple metastable states and a minimum energy pathway that does not parallel the axes of the variables used for the numerical integration make this system representative of difficulties commonly encountered in molecular simulations. Because the model is two-dimensional, we can readily visualize results and compare them with statistics independently computed using the grid-based scheme in Ref. 27.
@@
$X_{t+dt}=X_{t}-\nabla V(X_{t})dt+\sqrt{\frac{dt}{2\beta}}(Z_{t}+Z_{t-dt}),$ (35)
@@
Figure 1: Müller-Brown potential. Orange and red ovals indicate states $A$ and $B$, respectively.
@@ V.1 Comparison of algorithms
We measure convergence by computing the root mean square (RMS) difference in $\ln(\tilde{\pi}(x))$ from $-\beta V$, where $\tilde{\pi}(x)$ is an estimate of the steady-state distribution from the normalized histogram of samples. For the histogram, we use a uniform $50\times 50$ grid on the rectangle defined by the minimum and maximum values in the NEUS dataset. Since some grid regions are empty because they correspond to energies too high for NEUS to sample, we only compute errors over bins with $V<7$. We stop each simulation when the RMS difference drops below one. Results are shown in Figure 2. We see that WE requires about 73 times more iterations than NEUS, which in turn requires about 13 times more iterations than BAD-NEUS to reach the convergence criterion.
@@ V.3 Kinetic statistics
A common problem in molecular dynamics is that while equilibrium averages are relatively straightforward to calculate from biased simulations (via, for example, the various methods reviewed in Ref. 32), dynamical averages are much harder. In this section, we use BAD-NEUS to compute dynamical averages efficiently. To do so, we split the ensemble of transition paths based on the last metastable state visited as previously for NEUS [33, 16, 25]. We specifically compute the backward committor, $q_{-}$, which is the probability that the system last visited a reactant state $A$ rather than a product state $B$, and the transition path theory (TPT) rate, defined as the mean number of $A$ to $B$ transitions per unit time divided by the fraction of time spent last in $A$. The former can be obtained without saving additional information once we split the path ensemble. The forward committor, $q_{+}$, which is the probability that the system next visits state $B$ rather than state $A$, can also be obtained from NEUS but requires saving additional information [25]. Here, both of the systems that we consider are microscopically reversible, so we can obtain $q_{+}$ from $q_{+}=1-q_{-}$.
@@
where $\mathcal{T}_{t}$ is the operator that describes the evolution of expectations of functions:
$\mathcal{T}_{t}g(x,j)=E_{X_{0}=x,J_{0}=j}[f(X_{t},J_{t})].$ (40)
@@
Figure 6: Convergence of the TPT inverse rate estimate for the Müller-Brown system. The green line is the grid-based reference obtained from 44. For these simulations, there are 2000 walkers per stratum, $h=4$ iterations are retained, and the lag time is $\tau=10$ time steps.
@@
To compute reference values for these kinetic statistics, we use the grid-based approximation to the generator in Ref. 27, with the same grid parameters. Since the dynamics are microscopically reversible, we obtain the backward committor by solving for the forward committor using the approach in Ref. 27, then setting $q_{-}=1-q_{+}$. We solve for the reaction rate using
@@
Figure 6 shows estimates for the TPT rate. We see that BAD-NEUS converges several fold faster than NEUS. Each BAD-NEUS iteration requires an average of 407 time units of sampling (arising from 2000 walkers in each of 20 strata, with an average time before leaving the stratum of 10.2 time steps). Each iteration is therefore significantly less total computational cost than generating a single $A$ to $B$ transition on average (1200 time units, as evidenced by $k_{AB}^{-1}$) and is amenable to parallelization.
@@ VI Molecular Test System
All molecular dynamics simulations were performed with OpenMM [38]. We compute CVs including backbone RMSDs and fractions of native contacts using MDtraj [39]. To enable direct comparison with Ref. 34, we model NTL9(1-39) using the AMBER FF14SB force field [40] and the Hawkins, Cramer and Truhlar generalized Born implicit solvent model [41, 42]. We use a Langevin thermostat with a time step of 2 fs, a temperature of 300. K, and a friction constant of $\gamma=80$ ps-1, which corresponds to the high-viscosity case considered in Ref. 34.
@@
where $Q(X)$ is the fraction of native contacts using the definition of Ref. 43. We determine the native contacts from the crystal structure, PDB ID 2HBB[44]. We space 20 $Q_{k}$ values uniformly in the interval $[0.35,0.85]$ and set $\varepsilon_{k}=0.6(Q_{k+1}-Q_{k})$, so that the regions overlap. Thus there are 40 total strata. We define the folded state as full backbone RMSD $<0.28$ nm and $Q>0.85$; the denatured state has full backbone RMSD $>0.80$ nm and $Q<0.35$.
@@
Here we focus on the convergence of the folding rate; a detailed analysis of the NTL9 folding mechanism (with and without the C-terminal helix) based on potentials of mean force and committors will be presented elsewhere (see also Ref. 46). As mentioned above, we compare our results to those of Ref. 34, which estimates rates using haMSMs applied to data from WE (haMSM-WE). As discussed in the Introduction, BAD-NEUS goes beyond haMSM-WE by using the basis set (Markov model) to accelerate the convergence of the sampling rather than just the rate estimates.
Figure 10 shows the inverse rate constant obtained from BAD-NEUS and haMSM-WE [34]. The average inverse rate over the final 15 iterations of BAD NEUS is 4.74 ms, with a standard deviation of 0.99 ms, while the authors of Ref. 34 used a Bayesian bootstrapping approach to estimate a confidence interval of 0.17–1.9 ms, which they state is likely an underestimate of the true 95 percent confidence interval owing to the limited number of independent samples used in the analysis. Given that we have only a single BAD NEUS run, the standard deviation of the final 15 (highly correlated) inverse rate estimates is also likely a significant underestimate of our statistical error. We thus view the results in Figure 10 as in agreement.
@@
Both BAD-NEUS and haMSM-WE required significantly less aggregate simulation to compute the rate constant (an average) than the millisecond timescale of a single folding event. The haMSM-WE calculation used an aggregate simulation time of 252 $\mu$s, while we use an average of 4 $\mu$s per iteration, for a total of about 204 $\mu$s. The speedup is actually more than it appears because the haMSM-WE calculation is for the forward (folding) direction only; computing potentials of mean force and committors would also require the corresponding calculation for the backward (unfolding) direction. By contrast, the BAD-NEUS simulation provides these statistics in its present form.
Figure 9: Comparison of inverse rates for NTL9(1-39) folding from NEUS, BAD-NEUS and haMSM-WE. The haMSM-WE results are from Ref. 34 The shaded area shows the Bayesian $95\%$ credible interval reported in that publication.
@@ VII Conclusions
Trajectory stratification methods enable enhanced sampling for estimating both thermodynamic (equilibrium) and kinetic (nonequilibrium) statistics. Here, we introduced a new trajectory stratification method, BAD-NEUS, which converges faster than existing ones. We show that our method is a natural generalization of WE as originally formulated [1] but that, without modification, WE converges no faster than unbiased dynamics. The key modification that we introduce is the insertion of an approximation to the steady-state distribution before the resampling step. This approximation algorithm is designed so that it preserves the steady-state distribution of the dynamics, and therefore does not introduce systematic errors into the algorithm in the large data limit.
In the present study, we use a simple basis expansion to model the steady-state distribution, but our strategy is general, and alternatives are also possible. One can use neural networks to learn the steady-state distribution and/or the basis functions [26, 27], and one can incorporate memory into a basis expansion [47]. These alternatives, which can be used separately or together, alleviate the need to identify basis sets that describe the dynamics well and, in the case of memory, may reduce the sampling required. We thus expect our method to be a significant be step toward accurate estimates of rates for protein folding and similarly complex molecular conformational changes.
@@ Appendix A Derivation of (16)
Here we derive (16) by modifying the proof from Ref. 48. Let $Z_{t}$ be a Markov process, $\mathcal{T}_{t}$ be operator defined in (40), and $S_{n}=\sum_{t=0}^{n-1}g(Z_{t})$ be the partial sum of a function $g$. Let $\omega(z)$ satisfy the linear problem
@@
We refer the reader to Ref. 48 for a more technical discussion and conditions under which the relevant expectations are bounded such that we may apply the optional stopping theorem. Then, for any stopping time $T$ with $\operatorname{\mathbbm{E}}[T]<\infty$,

### v2

@@ V Two-dimensional test system
$V_{\rm MB}(u,v)=\frac{1}{20}\sum_{i=1}^{4}C_{i}\exp[a_{i}(u-v_{i})^{2}\\ +b_{i}(u-u_{i})(v-v_{i})+c_{i}(v-v_{i})^{2}].$ (35)
For all results shown, we use $C_{i}=\{-200,-100,-170,15\}$, $a_{i}=\{-1,-1,-6.5,0.7\}$, $b_{i}=\{0,0,11,0.6\}$, $c_{i}=\{-10,-10,-6.5,0.7\}$, $u_{i}=\{1,-0.27,-0.5,-1\}$, $v_{i}=\{0,0.5,1.5,1\}$. The potential with these parameter choices is shown in Figure 1. The presence of multiple metastable states and a minimum energy pathway that does not parallel the axes of the variables used for the numerical integration make this system representative of difficulties commonly encountered in molecular simulations. Because the model is two-dimensional, we can readily visualize results and compare them with statistics independently computed using the grid-based scheme in Ref. 30.
@@
$X_{t+dt}=X_{t}-\nabla V_{\rm MB}(X_{t})dt+\sqrt{\frac{dt}{2\beta}}(Z_{t}+Z_{t-dt}),$ (36)
@@
Figure 1: Müller-Brown potential. Orange and red ovals indicate states $A$ and $B$, respectively. Contours are spaced every $k_{B}T$
@@ V.1 Comparison of algorithms
We measure convergence by computing the root mean square (RMS) difference in $\ln(\tilde{\pi}(x))$ from $-\beta V_{\rm MB}(x)$, where $\tilde{\pi}(x)$ is an estimate of the steady-state distribution from the normalized histogram of samples. For the histogram, we use a uniform $50\times 50$ grid on the rectangle defined by the minimum and maximum values in the NEUS dataset. Since some grid regions are empty because they correspond to energies too high for NEUS to sample, we only compute errors over bins with $V_{\rm MB}<7$. We stop each simulation when the RMS difference drops below one. Results are shown in Figure 2. We see that WE (without any acceleration strategy) requires about 73 times more iterations than NEUS, which in turn requires about 13 times more iterations than BAD-NEUS to reach the convergence criterion.
@@ V.3 Kinetic statistics
A common problem in molecular dynamics is that while equilibrium averages are relatively straightforward to calculate from biased simulations (via, for example, the various methods reviewed in Ref. 35), dynamical averages are much harder. In this section, we use BAD-NEUS to compute dynamical averages efficiently. To do so, we split the ensemble of transition paths based on the last metastable state visited as previously for NEUS [36, 17, 26]. We specifically compute the backward committor, $q_{-}$, which is the probability that the system last visited a reactant state $A$ rather than a product state $B$, and the transition path theory (TPT) rate, defined as the mean number of $A$ to $B$ transitions per unit time divided by the fraction of time spent last in $A$. The former can be obtained without saving additional information once we split the path ensemble. The forward committor, $q_{+}$, which is the probability that the system next visits state $B$ rather than state $A$, can also be obtained from NEUS but requires saving additional information [26]. Here, both of the systems that we consider are microscopically reversible, so we can obtain $q_{+}$ from $q_{+}=1-q_{-}$.
@@
where $B^{c}$ is the complement of $B$, and $\mathcal{T}_{t}$ is the operator that describes the evolution of expectations of functions:
$\mathcal{T}_{t}g(x,j)=E_{X_{0}=x,J_{0}=j}[g(X_{t},J_{t})].$ (41)
@@
To compute reference values for these kinetic statistics, we use the grid-based approximation to the generator in Ref. 30, with the same grid parameters. Since the dynamics are microscopically reversible, we obtain the backward committor by solving for the forward committor using the approach in Ref. 30, then setting $q_{-}=1-q_{+}$. We solve for the reaction rate using
@@
Figure 6 shows estimates for the TPT rate. We see that BAD-NEUS converges several fold faster than NEUS, Each BAD-NEUS iteration requires an average of 407 time units of sampling (arising from 2000 walkers in each of 20 strata, with an average time before leaving the stratum of 10.2 time steps). Each iteration is therefore significantly less total computational cost than generating a single $A$ to $B$ transition on average (1200 time units, as evidenced by $k_{AB}^{-1}$) and is amenable to parallelization.
@@
Figure 6: Convergence of the TPT inverse rate estimate for the Müller-Brown system. The green line is the grid-based reference obtained from 45. For these simulations, there are 2000 walkers per stratum, $h=4$ iterations are retained, and the lag time is $\tau=10$ time steps.
@@ VI Molecular Test System
All molecular dynamics simulations were performed with OpenMM [41]. We compute CVs including backbone RMSDs and fractions of native contacts using MDtraj [42]. To enable direct comparison with Ref. 37, we model NTL9(1-39) using the AMBER FF14SB force field [43] and the Hawkins, Cramer and Truhlar generalized Born implicit solvent model [44, 45]. We use a Langevin thermostat with a time step of 2 fs, a temperature of 300. K, and a friction constant of $\gamma=80$ ps-1, which corresponds to the high-viscosity case considered in Ref. 37.
@@
where $Q(X)$ is the fraction of native contacts using the definition of Ref. 46. We determine the native contacts from the crystal structure, PDB ID 2HBB [47]. We space 20 $Q_{k}$ values uniformly in the interval $[0.35,0.85]$ and set $\varepsilon_{k}=0.6(Q_{k+1}-Q_{k})$, so that the regions overlap. Thus there are 40 total strata. We define the folded state as full backbone RMSD $<0.28$ nm and $Q>0.85$; the denatured state has full backbone RMSD $>0.80$ nm and $Q<0.35$.
@@
Here we focus on the convergence of the folding rate; a detailed analysis of the NTL9 folding mechanism (with and without the C-terminal helix) based on potentials of mean force and committors will be presented elsewhere (see also Ref. 49). As mentioned above, we compare our results to those of Ref. 37, which estimates rates using haMSMs applied to data from WE (haMSM-WE). As discussed in the Introduction, BAD-NEUS goes beyond haMSM-WE by using the basis set (Markov model) to accelerate the convergence of the sampling rather than just the rate estimates.
Figure 9 shows the inverse rate constant obtained from BAD-NEUS and haMSM-WE [37]. The average inverse rate over the final 15 iterations of BAD NEUS is 4.74 ms, with a standard deviation of 0.99 ms, while the authors of Ref. 37 used a Bayesian bootstrapping approach to estimate a confidence interval of 0.17–1.9 ms, which they state is likely an underestimate of the true 95 percent confidence interval owing to the limited number of independent samples used in the analysis. Given that we have only a single BAD NEUS run, the standard deviation of the final 15 (highly correlated) inverse rate estimates is also likely a significant underestimate of our statistical error. We thus view the results in Figure 9 as in agreement.
@@
We run each walker until its index process changes and then an additional 10 ps, so the trajectory length and time per iteration are random variables. For the NTL9 simulations, the average trajectory length was 50 ps, and the average time per iteration was 4 $\mu$s. The results in Figure 9 thus required about 204 $\mu$s in total, comparable to the 252 $\mu$s for the haMSM-WE calculation. The time to compute the rate constant by BAD-NEUS and haMSM-WE is much less than the millisecond timescale of a single folding event, which we also expect to be the time to converge WE without any acceleration strategy. BAD-NEUS actually provides some speedup relative to haMSM-WE in the sense that it provides potentials of mean force and committors in its present form, whereas a calculation in the unfolding direction (performed either separately or simultaneously, as here) would be required to obtain these statistics from haMSM-WE. By the same token, we likely could reduce the computation time for BAD-NEUS by “teleporting” walkers from $B$ to $A$ and computing only the forward rate [50]. However, we expect most investigators would seek the mechanistic information provided by the potentials of mean force and committor (which can be used to also obtain reactive currents [24]) if they are investing the computational resources to compute the rate, and free energies computed from ratios of forward and backward rates can serve as checks on the rates [51, 52].
Figure 9: Comparison of inverse rates for NTL9(1-39) folding from BAD-NEUS and haMSM-WE. The haMSM-WE results are from Ref. 37 The shaded area shows the Bayesian $95\%$ credible interval reported in that publication.
@@ VII Conclusions
Trajectory stratification methods enable enhanced sampling for estimating both thermodynamic (equilibrium) and kinetic (nonequilibrium) statistics. Here, we introduced a new trajectory stratification method, BAD-NEUS, which converges faster than existing ones. We show that our method is a natural generalization of NEUS and EM, which in turn can be formulated as extensions of WE. In the process, we show that WE as originally formulated [1] converges no faster than unbiased dynamics. The key modification that we introduce is the insertion of an approximation to the steady-state distribution before the resampling step. Importantly, we design this approximation algorithm so that it preserves the steady-state distribution of the dynamics, and therefore does not introduce systematic errors into the algorithm in the large-data limit.
Our approach defines precise ingredients needed for iterative restart strategies to converge to the correct fixed point (i.e., a fixed point consistent with unbiased dynamics). In this way, it both restricts and generalizes iterative restart strategies previously proposed [13, 8], and we provide numerical demonstrations that such strategies can accelerate convergence of both equilibrium and kinetic statistics. In the future, it would be interesting to compare BAD-NEUS with the WE milestoning approach, which instead embeds WE simulations within coarse regions (i.e., between milestones) [53, 54]. While this approach also combines fine and coarse information, it is fundamentally different from BAD-NEUS and haMSM-WE in that it introduces an additional sampling rather than analysis algorithm at the fine level. Indeed, adding an element of stratification to the WE simulations within WE milestoning may result in further speedups. The similarities and differences between various approaches highlights the importance of a theoretical framework for relating algorithms like the one that we introduce here; we believe it can serve as the basis for a thorough, systematic cataloging of methods for computing dynamical statistics, analogous to recent attempts for enhanced sampling methods for equilibrium statistics [35].
In the present study, we use a simple basis expansion to model the steady-state distribution, but our strategy is general, and alternatives are also possible. One could use neural networks to learn the steady-state distribution and/or the basis functions [29, 30]. One could also incorporate memory [55]. These alternatives, which could be used separately or together, alleviate the need to identify basis sets and/or CVs that describe the dynamics well and, in the case of memory, may reduce the sampling required. We thus expect our method to be a significant step toward accurate estimates of rates for protein folding and similarly complex molecular conformational changes.
@@ Appendix A Derivation of (16)
Here we derive (16) by modifying the proof from Ref. 56. Let $Z_{t}$ be a Markov process, $\mathcal{T}_{t}$ be operator defined in (41), and $S_{n}=\sum_{t=0}^{n-1}g(Z_{t})$ be the partial sum of a function $g$. Let $\omega(z)$ satisfy the linear problem
@@
We refer the reader to Ref. 56 for a more technical discussion and conditions under which the relevant expectations are bounded such that we may apply the optional stopping theorem. Then, for any stopping time $T$ with $\operatorname{\mathbbm{E}}[T]<\infty$,